# CMB delensing with deep learning

### v2

@@ Abstract
The cosmic microwave background (CMB) stands as a pivotal source for studying weak gravitational lensing. While the lensed CMB aids in constraining cosmological parameters, it simultaneously smooths the original CMB’s features. The angular power spectrum of the unlensed CMB showcases sharper acoustic peaks and more pronounced damping tails, enhancing the precision of inferring cosmological parameters that influence these aspects. Although delensing diminishes the $B$-mode power spectrum, it facilitates the pursuit of primordial gravitational waves and enables a lower variance reconstruction of lensing and additional sources of secondary CMB anisotropies. In this work, we explore the potential of deep learning techniques, specifically the U-Net++ algorithm, to play a pivotal role in CMB delensing. We analyze three fields, namely $T$, $Q$, and $U$ sky maps, present the angular power spectra of the CMB delensed $TT$, $EE$, $BB$, and $TE$, and compare them with the unlensed CMB angular power spectra. Our findings reveal that the angular power spectrum of the lensed CMB, processed by U-Net++, closely aligns with that of the unlensed CMB. Thus, U-Net++ based CMB delensing proves to be effective in mitigating the impacts of weak gravitational lensing, paving the way for enhancing the CMB delensing power spectrum in forthcoming CMB experiments. The code utilized for this analysis is available on GitHub.
keywords:
cosmology: cosmic microwave background – cosmology: gravitational lensing
pubyear: 2015pagerange: CMB delensing with deep learningCMB delensing with deep learning
1 Introduction
@@
The anisotropies of the cosmic microwave background (CMB) offer crucial insights into the early universe. Observations of such anisotropies have made significant contributions to establishing the current standard model of cosmology. In the last few decades, three generations of satellite experiments, namely COBE (Smoot et al., 1992; Fixsen et al., 1996; Bennett et al., 1996), WMAP (Spergel et al., 2003; Komatsu et al., 2009, 2011), and Planck (Ade et al., 2014b; Ade et al., 2016; Aghanim et al., 2020), as well as numerous ground-based experiments (e.g., DASI (Leitch et al., 2002), CBI (Readhead et al., 2004), SPTpol (Carlstrom et al., 2011), BICEP (Barkats et al., 2014; Ade et al., 2014c), Keck Array (Staniszewski et al., 2012), ACT (Coulton et al., 2023), and Ali-CPT (Li et al., 2019)) and balloon experiments (e.g., BOOMERanG (Masi et al., 2002), EBEX (Aboobaker et al., 2018), SPIDER (Crill et al., 2008)), have been carried out to precisely measure the temperature and polarization power spectrum of the CMB. Moreover, there are several ongoing plans for diverse experimental projects, including LiteBIRD (Allys et al., 2023), CMB-S4 (Abazajian et al., 2016, 2022), and PIPER (Lazear et al., 2014), being devoted to the detection of CMB polarization signals.
@@
The generation of the stochastic background of gravitational waves, known as primordial gravitational waves (PGWs), is a fundamental prediction in any cosmological inflation model (Lyth & Riotto, 1999; Baumann, 2011). The characteristic of this signal encodes unique information about the physics of the early universe and its subsequent evolution, providing an exciting and powerful window into the universe’s origin and evolution.
@@
The tensor-to-scalar ratio $r$ parameterizes the amplitude of PGWs and connects to the energy scale at which inflation occurred. Therefore, the detection of PGWs is expected to reveal the physical properties of the universe’s early stages. Fortunately, these PGWs have distinct imprints in the polarized anisotropy of the CMB, displaying a spiral-like B-mode pattern that enables the extraction of the PGW signal from the polarized B-mode of the CMB (Ade et al., 2014a; Abazajian et al., 2016). The detection of B-mode signals presents an exceptionally challenging task, particularly due to the uncertain extent of foreground contamination and the mixing of relatively strong E-mode signals with B-mode signals induced by weak gravitational lensing (Ade et al., 2014a). Therefore, the primary scientific objectives of CMB polarization observations currently entail determining the extent of foreground contamination and precisely measuring the effect of gravitational lensing (Ade et al., 2019). Our research primarily centers on the gravitational lensing phenomenon associated with CMB.
@@
CMB lensing has been extensively studied in theory (Lewis & Challinor, 2006; Okamoto & Hu, 2003; Bartelmann & Schneider, 2001). After decoupling from matter at the last scattering surface, CMB photons propagate freely and are gravitationally deflected by the universe’s large scale distribution of matter. The lensing effect results in subtle imprints on the CMB temperature and polarization anisotropies. These can be utilized to create a map of the lensing potential that determines the lensing deflections’ gradient (Hu, 2000). Weak gravitational lensing smoothes the acoustic peaks of the CMB angular power spectra, transferring power from larger angular scales to smaller scales, and converting E-mode polarization to B-mode polarization (Zaldarriaga & Seljak, 1998; Dodelson & Schmidt, 2021). In addition, weak gravitational lensing can induce small distortions in the CMB, and these distortions can be detected by modifying its primordial morphology in an anisotropic form.
@@
Given that the anisotropy during the final scattering can be approximated as Gaussian, the non-Gaussian structures observed in the lensed CMB sky provide additional information about the properties of gravitational lensing material. This is crucial for a further understanding of important details regarding the distribution of matter (Lewis et al., 2000; Lewis & Challinor, 2006; Lewis, 2013). Therefore, weak gravitational lensing of the CMB both promotes and hinders our comprehension of the history and content of the universe.
@@
There are several advantages associated with mitigating the impacts of lensing on the observed CMB temperature and polarization maps (Lewis, 2005; Lewis & Challinor, 2006; Lewis, 2013; Dodelson & Schmidt, 2021). Firstly, the average temperature of the CMB remains unchanged by lensing as the gravitational effect only adjusts arrival directions and not the surface brightness (Lewis, 2005, 2013). Secondly, the distortion of light paths traveling from far sources to reach us is caused by the gravitational effect of universe’s inhomogeneities. The reason lensing is so promising is that it enables probing all clustering stress-energy components in the universe through spacetime perturbations since light paths react to mass. Measurement of these distortions provides insight into the mass distribution of the universe (Lewis, 2005; Lewis & Challinor, 2006; Lewis, 2013).
@@
However, gravitational lensing has the potential to smoothen acoustic peaks in the CMB. This can be comprehended by considering that when photons get deflected, features of a fixed angular size can either be magnified or de-magnified, causing sharp features in the power spectrum to be blurred across a range of scales. The angular scale of sharp features in the power spectrum is easier to measure than broad humps; consequently, gravitational lensing weakens our ability to precisely measure acoustic peak positions in the CMB power spectra (Lewis, 2005; Lewis & Challinor, 2006). The temperature power spectrum undergoes lensing by $0.2\%$ at $\ell\sim 2000$, but smaller scales experience changes at the percentage level. The impact on the B-mode polarization power spectrum is more significant, with power increase of $6\%$ across all scales (Lewis & Challinor, 2006).
@@
Delensing reverses this peak smoothing, providing sharper peaks with a more precisely measurable angular scale. Similar observations can be made about the measurement of peak heights. Weak gravitational lensing alters the CMB power spectra, induces non-Gaussian, and generates a B-mode polarization signal, which causes confusion for the signal from PGWs (Lewis, 2005; Lewis & Challinor, 2006).
@@
In recent years, deep learning and artificial intelligence techniques have developed rapidly, gaining widespread attention and significant application in various disciplines. Astronomy research has been catching up, and many studies have emerged, applying deep learning for data analysis (Gupta et al., 2018; Caldeira et al., 2019; Zhou et al., 2018; Zhang et al., 2019; Zhou et al., 2019; Springer et al., 2020; Makinen et al., 2021; Guzman & Meyers, 2022; Ni et al., 2022; Gao et al., 2023). These studies have demonstrated the high effectiveness of deep learning in image reconstruction and segmentation tasks, enabling the detection of features at the pixel level (Wang et al., 2018; Ghosh et al., 2019; Minaee et al., 2021).
@@
This work adopts the sky map segmentation method proposed by Makinen et al. (2021). The method relies on the Healpix pixelization scheme, which divides the two-dimensional spherical sky map into multiple two-dimensional plane images. We apply this sky map segmentation method to attempt a novel CMB delensing approach that can remove the lensing effect.
The paper is organized as follows. In Section 2, we present a fundamental overview of CMB and lensing, along with a clear problem statement. This section also encompasses simulations of both CMB lensing and unlensing data, as well as noise simulation. Section 3 delves into the details of the U-Net++ architecture, data pre-processing techniques, and the predictions generated by U-Net++. In Section 4, we provide a comprehensive analysis of the obtained results. Finally, Section 5 provides the concluding remarks.
2 Data Simulation
@@
In this section, we will provide a comprehensive description of the simulated CMB and noisy sky maps used in our analysis. This includes detailed information on both the lensed temperature and polarization maps with noise, as well as the unlensed temperature and polarization maps.
@@
For our simulations, we employed the publicly available package (Lenspyx [^1] https://github.com/carronj/lenspyx(Lewis, 2005; Diego-Palazuelos et al., 2020), a Python package specifically designed for simulating lensed CMB maps on a curved sky. We also relied on three additional software packages, namely CAMB [^2] https://camb.info/ (Lewis et al., 2000), LensPix [^3] https://cosmologist.info/lenspix/ (Lewis, 2005) and Healpy [^4] https://github.com/healpy/healpy/ (Górski et al., 2005; Zonca et al., 2019), as the foundational components for developing detailed curved-full-sky simulations of both lensed and unlensed CMB. By incorporating the Planck 2018 CMB lensing pipeline (plancklens), Lenspyx has the ability to replicate both the published map and band-powers.
@@
where, $\mathcal{M}$ represents the full sky map, and $\mathcal{F}\in\{T,Q,U\}$. Smooth represents the telescope’s beam, and we have chosen a simple Gaussian beam. Lenspyx is an abbreviation for the process of converting unlensed maps into lensed maps. $N$ represents the noise component in the image. Here, Smooth and Lenspyx can be regarded as operators, corresponding to the operations of achieving smoothing and lensing, respectively.
During the simulation process, we employed the HEALPix pixelization strategy. Given the significant importance of small-scale analysis, we deliberately selected a sky map with a resolution of $N_{side}=2048$, corresponding to an angular resolution of $1.718$ arcminutes.
We have considered the simple instrumental effect, namely the influence of the Gaussian beam. In order to more accurately simulate the observational data, we adopt the full width at half maximum (FWHM) from existing full sky observation experiments in subsequent simulations to convolve the sky map.
2.1 Noise simulation
The noise in CMB detectors is often approximated as Gaussian white noise, and its angular power spectral density can be defined in the following (Hanany et al., 2019; Diego-Palazuelos et al., 2020; Wolz et al., 2023):
@@ 2.1 Noise simulation
where $X\in\{T,E,B\}$, $\theta_{\text{FWHM}}$ stands for the FWHM of the telescope, which is related to the observation frequency. Additionally, $s$ denotes the telescope’s sensitivity, which is defined as (Wu et al., 2014; Ade et al., 2015; Wu et al., 2020)
$s^{-1}[\mu\text{K}~{}\text{rad}]=\frac{\text{NET}[\mu\text{K}~{}\sqrt{\text{sec}}]~{}\sqrt{4\pi f_{\text{sky}}[\text{rad}^{2}]}}{\sqrt{N_{\text{det}}~{}Y~{}\Delta T[\text{sec}]}},$ (3)
where NET represents the instrument equivalent temperature, $N_{\text{det}}$ is the number of detectors, $f_{\text{sky}}$ denotes the observed sky fraction, and we assume an ideal case where yield $Y=1$, and $\Delta T$ represents the observation duration.
Our research primarily aims to comprehensively examine the potential of delensing rather than extensively investigate the specific impact of noise in the delensing process. To fulfill this research objective, we have intentionally selected a specific frequency, namely $75$ GHz (Akrami et al., 2020; Hanany et al., 2019; Allys et al., 2023), as the sole focus of our in-depth analysis. Furthermore, considering that our analysis relies on full-sky data, we employed the instrument parameters of PICO (Hanany et al., 2019), the next-generation space telescope, to simulate the noise, as detailed in Table 1.
Table 1: Instrument and noise specifications used to produce the simulations in this work. In the table, $\theta_{\text{FWHM}}$ represents the telescope’s full width at half maximum; NET($T/P$) denotes the effective temperature of the instrument, with $T$ referring to the temperature field and $P$ indicating the polarization field; $N_{\text{det}}$ stands for the number of detectors; $\Delta T$ signifies the total observation time; $f_{\rm sky}$ represents the fraction of the sky observed within the $\Delta T$ timeframe; and $Y$ assumes a value of $1$ under ideal conditions.
[…]
$10.7$ $38/2.1$ $1020$ $80$ $1$ $1.5\times 10^{5}$
After substituting the parameters into Equaitons (2) and (3), we can calculate the temperature and polarization angular power spectrua of noise, as illustrated in Figures 1 and 2. For the noise of each component, we generate corresponding full sky maps based on their angular power spectra and stack these maps onto the sky maps with lensing effects. Through this process, we obtain the training dataset required for the U-Net++ model in subsequent studies.
2.2 The CMB temperature and lensing potential
The CMB radiation field is represented by the temperature anisotropy, denoted as $T(\hat{n})$, and the polarization, denoted as $P(\hat{n})$, in the spatial direction $\hat{n}$ of the celestial sphere. We observe CMB temperature changes projected on a 2D spherical surface sky, and it is now habitual in the literature to expand the temperature field using spherical harmonics.
The temperature fluctuation of the CMB on the spherical surface is described by a scalar field, consisting of small fluctuations $\Delta T(\hat{n})$ at a level of $10^{-5}$ relative to the average value $T_{0}=2.725$ K. For full-sky observations, the temperature field T can be expressed through the spherical harmonic decomposition using spin-0 spherical harmonics $Y_{lm}(\hat{n})$,
$\Delta T(\hat{n})=\sum_{\ell=0}^{l_{max}}\sum_{m=-\ell}^{m=\ell}a_{\ell m}Y_{\ell m}(\hat{n}).$ (4)
All the information contained in the temperature field $T(\hat{n})$ is included in the space-time dependent amplitudes $a_{\ell m}$, where $a_{\ell m}$ represents the spherical harmonic coefficient and can be expressed using the following formula,
@@ 2.2 The CMB temperature and lensing potential
$<a_{\ell m}a_{\ell^{\prime}m^{\prime}}^{*}>=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}C_{\ell},$ (6)
@@
We can simulate the angular power spectrum $C_{\ell}$ with CAMB or CLASS [^5] https://lesgourg.github.io/class_public/class.html (Lucca et al., 2020; Di Dio et al., 2013), as shown in Fig. 1. The figure demonstrates that as the scale decreases, the difference between lensed and unlensed data amplifies. This phenomenon occurs due to the deflection imparted on CMB photons by each encountered potential, resulting in an accumulation of effects on the CMB power spectrum at small scales. Consequently, the lensing effect distorts the original power spectrum and introduces non-Gaussianity in the lensed CMB (Lewis & Challinor, 2006). It should be noted that non-linear evolution also contributes to an increase in power on smaller scales.
Weak lensing of the CMB deflects photons coming from an original direction $\hat{n}^{\prime}$ on the last scattering surface to direction $\hat{n}$ on the observed sky, so a lensed CMB temperature field, $T(\theta,\phi)$, is given by $\widetilde{X}(\hat{n})=X(\hat{n}^{\prime})$ in terms of the unlensed field $X=T$ (Lewis, 2005). Thus the position in the sky where we finally see the CMB photons is determined by the integral of the gravitational potential along the line of sight to the last scattering surface.
We propose the definition of an integrated lensing potential, denoted as $\psi$. The deflection vector is expressed as the gradient of the lensing potential, $\nabla\psi(\hat{n})$, where $\nabla$ represents the covariant derivative on the sphere. The vector $\hat{n}^{\prime}$ is derived from $\hat{n}$ by moving its geodesic at one end of the surface of the unit sphere along the $\hat{n}^{\prime}$ direction by a distance $\nabla\psi(\hat{n})$. Then the unlensed CMB photon with direction $\hat{n}$ becomes a lensed CMB photon with direction $\hat{n}=\hat{n}^{\prime}+\nabla\psi(\hat{n})$, through the weak gravitational lensing. Thus, the lensed CMB temperature can be written as follows,
$\widetilde{T}(\hat{n})=T(\hat{n}^{\prime})=T(\hat{n}+\nabla\psi(\hat{n})),$ (8)
where lensing potential $\psi(\hat{n})$ can be defined as (Lewis & Challinor, 2006; Hassani et al., 2016),
@@
where $T(k,\eta)$ is the appropriate transfer function, $P_{R}(k)$ is the primordial power spectrum. We will employ either CAMB or CLASS to simulate the power spectrum of the lensing potential.
The CMB photons are deflected by the gravitational lensing effect of massive cosmic structures as they travel across the universe. Researchers have been able to measure this gravitational lensing of the CMB across the full sky using data from CMB satellites. Figure 1 shows the lensed TT angular power spectrum and the unlensed TT power spectrum, along with their ratios. On large scale, the anisotropies of the CMB are primarily dominated by emissions from the last scattering surface at a redshift of $z\sim 1100$. However, on smaller scale, the CMB is more significantly influenced by what are known as secondary effects. These secondary anisotropies arise as a result of interactions between CMB photons and matter along the line of sight. Consequently, as seen in the bottom panel of Figure 1, the small-scale angular power spectrum is more significantly affected by the lensing potential.
Figure 1: Effect of lensing on the CMB temperature power spectrum. Top panel: The solid green line represents the unlensed CMB TT angular power spectrum. The light green dashed line represents the lensed CMB TT angular power spectrum. The gray dashed line depicts the contribution of noise to the TT angular power spectrum. Bottom panel: The fractional difference between these very similar spectra. The red solid line illustrates the ratio between the unlensed TT spectrum and the lensed TT spectrum, while the gray solid line represents the reference line where the ratio equals $1$.

### v3

@@ Abstract
The CMB stands as a pivotal source for studying weak gravitational lensing. While the lensed CMB aids in constraining cosmological parameters, it simultaneously smooths the original CMB’s features. The angular power spectrum of the unlensed CMB showcases sharper acoustic peaks and more pronounced damping tails, enhancing the precision of inferring cosmological parameters that influence these aspects. Although delensing diminishes the B-mode power spectrum, it facilitates the pursuit of primordial gravitational waves and enables a lower variance reconstruction of lensing and additional sources of secondary CMB anisotropies. We employed the UNet++ algorithm to perform operations and analysis on CMB delensing, presenting the angular power spectra of TT, EE, and BB after CMB delensing, and compared them with those obtained using the QE delensing algorithm. The lensing CMB sky map and full-sky angular power spectrum processed by the UNet++ algorithm are very close to those of the CMB without lensing effects, and the error is more than 10 times smaller than that given by the QE algorithm. The code utilized for this analysis is publicly available.
I Introduction
@@
The anisotropies of the cosmic microwave background (CMB) offer crucial insights into the early universe. Observations of such anisotropies have made significant contributions to establishing the current standard model of cosmology. In the last few decades, three generations of satellite experiments, namely COBE COBE:1992syq; Fixsen:1996nj; Bennett:1996ce, WMAP WMAP:2003elm; WMAP:2008lyn; WMAP:2010qai, and Planck Planck:2013pxb; Planck:2015fie; Planck:2018vyg, as well as numerous ground-based experiments (e.g., DASI Leitch:2002fe, CBI Readhead:2004xg, SPTpol Carlstrom:2009um, BICEP BICEP1:2013sbv; BICEP2:2014dgt, Keck Array Staniszewski:2012eqo, ACT ACT:2023wcq, and Ali-CPT Li:2017drr) and balloon experiments (e.g., BOOMERanG Masi:2002hp, EBEX EBEX:2017etk, SPIDER Crill:2008rd), have been carried out to precisely measure the temperature and polarization power spectrum of the CMB. Moreover, there are several ongoing plans for diverse experimental projects, including LiteBIRD LiteBIRD:2022cnt, CMB-S4 CMB-S4:2016ple; CMB-S4:2022ght, AliCPT Li:2017drr, and PIPER Lazear:2014bga, being devoted to the detection of CMB polarization signals.
@@
The generation of the stochastic background of gravitational waves, known as primordial gravitational waves (PGWs), is a fundamental prediction in any cosmological inflation model Lyth:1998xn; Baumann:2009ds. The characteristic of this signal encodes unique information about the physics of the early universe and its subsequent evolution, providing an exciting and powerful window into the universe’s origin and evolution.
@@
The tensor-to-scalar ratio $r$ parameterizes the amplitude of PGWs and connects to the energy scale at which inflation occurred. Therefore, the detection of PGWs is expected to reveal the physical properties of the universe’s early stages. Fortunately, these PGWs have distinct imprints in the polarized anisotropy of the CMB, displaying a spiral-like B-mode pattern that enables the extraction of the PGW signal from the polarized B-mode of the CMB BICEP2:2014owc; CMB-S4:2016ple. The detection of B-mode signals presents an exceptionally challenging task, particularly due to the uncertain extent of foreground contamination and the mixing of relatively strong E-mode signals with B-mode signals induced by weak gravitational lensing BICEP2:2014owc. Therefore, the primary scientific objectives of CMB polarization observations currently entail determining the extent of foreground contamination and precisely measuring the effect of gravitational lensing SimonsObservatory:2018koc. Our research primarily centers on the gravitational lensing phenomenon associated with CMB.
@@
CMB lensing has been extensively studied in theory Lewis:2006fu; Okamoto:2003zw; Bartelmann:1999yn. After decoupling from matter at the last scattering surface, CMB photons propagate freely and are gravitationally deflected by the universe’s large scale distribution of matter. The lensing effect results in subtle imprints on the CMB temperature and polarization anisotropies. These can be utilized to create a map of the lensing potential that determines the lensing deflections’ gradient Hu:2000ee. Weak gravitational lensing smoothes the acoustic peaks of the CMB angular power spectra, transferring power from larger angular scales to smaller scales, and converting E-mode polarization to B-mode polarization Zaldarriaga:1998ar; DODELSON2021373. In addition, weak gravitational lensing can induce small distortions in the CMB, and these distortions can be detected by modifying its primordial morphology in an anisotropic form.
@@
Given that the anisotropy during the final scattering can be approximated as Gaussian, the non-Gaussian structures observed in the lensed CMB sky provide additional information about the properties of gravitational lensing material. This is crucial for a further understanding of important details regarding the distribution of matter Lewis:1999bs; Lewis:2006fu; Lewis:2013hha. Therefore, weak gravitational lensing of the CMB both promotes and hinders our comprehension of the history and content of the universe.
@@
There are several advantages associated with mitigating the impacts of lensing on the observed CMB temperature and polarization maps Lewis:2005tp; Lewis:2006fu; Lewis:2013hha; DODELSON2021373. Firstly, the average temperature of the CMB remains unchanged by lensing as the gravitational effect only adjusts arrival directions and not the surface brightness Lewis:2005tp; Lewis:2013hha. Secondly, the distortion of light paths traveling from far sources to reach us is caused by the gravitational effect of universe’s inhomogeneities. The reason lensing is so promising is that it enables probing all clustering stress-energy components in the universe through spacetime perturbations since light paths react to mass. Measurement of these distortions provides insight into the mass distribution of the universe Lewis:2005tp; Lewis:2006fu; Lewis:2013hha.
@@
However, gravitational lensing has the potential to smoothen acoustic peaks in the CMB. This can be comprehended by considering that when photons get deflected, features of a fixed angular size can either be magnified or de-magnified, causing sharp features in the power spectrum to be blurred across a range of scales. The angular scale of sharp features in the power spectrum is easier to measure than broad humps; consequently, gravitational lensing weakens our ability to precisely measure acoustic peak positions in the CMB power spectra Lewis:2005tp; Lewis:2006fu. The temperature power spectrum undergoes lensing by $0.2\%$ at $\ell\sim 2000$, but smaller scales experience changes at the percentage level. The impact on the B-mode polarization power spectrum is more significant, with power increase of $6\%$ across all scales Lewis:2006fu.
@@
Delensing reverses this peak smoothing, providing sharper peaks with a more precisely measurable angular scale. Similar observations can be made about the measurement of peak heights. Weak gravitational lensing alters the CMB power spectra, induces non-Gaussian, and generates a B-mode polarization signal, which causes confusion for the signal from PGWs Lewis:2005tp; Lewis:2006fu.
@@
In recent years, deep learning and artificial intelligence techniques have developed rapidly, gaining widespread attention and significant application in various disciplines. Astronomy research has been catching up, and many studies have emerged, applying deep learning for data analysis Gupta:2018eev; Caldeira:2018ojb; zhou2018unet++; Zhang:2019ryt; zhou2019unet++; Springer:2018aak; Makinen:2020gvh; Guzman:2021ygf; Ni:2022kxn; Gao:2022xdb; Yan:2024czw. These studies have demonstrated the high effectiveness of deep learning in image reconstruction and segmentation tasks, enabling the detection of features at the pixel level wang2018interactive; ghosh2019understanding; minaee2021image.
@@
This work adopts the sky map segmentation method proposed by Makinen Makinen:2020gvh. The method relies on the HEALPix pixelization scheme, which divides the two-dimensional spherical sky map into multiple two-dimensional plane images. We apply this sky map segmentation method to attempt a novel CMB delensing approach that can remove the lensing effect.
The paper is organized as follows. In Section II, we present the simulation of CMB and the preprocessing of simulated data. In Section III, we describe the methods of quadratic estimator (QE) delensing and UNet++ delensing. In Section IV, we provide a comprehensive analysis of the obtained results. Finally, Section V provides the concluding remarks.
II Data Simulation and Pre-processing
@@
In this section, we will describe the simulated CMB and noisy sky maps used for our analysis. This includes detailed information on lensed temperature and polarization maps with noise and instrumental effects, as well as unlensed temperature and polarization maps.
@@
For our simulations, we employed the publicly available package, Lenspyx Lewis:2005tp; Diego-Palazuelos:2020lme, a Python package specifically designed for simulating lensed CMB maps on a curved sky. We also relied on three additional software packages, namely CAMB Lewis:1999bs, LensPix Lewis:2005tp and Healpy Gorski:2004by; Zonca:2019vzt, as the foundational components for developing detailed curved-full-sky simulations of both lensed and unlensed CMB. By incorporating the Planck 2018 CMB lensing pipeline (plancklens), Lenspyx has the ability to replicate both the published map and band-powers.
@@
$\mathcal{M}^{\mathcal{F}}_{{\rm lensed}}={\rm Smooth}[{\rm Lenspyx}(\mathcal{M}^{\mathcal{F}}_{{\rm unlensed}})+\mathcal{M}^{\mathcal{F}}_{\rm N}],$ (1)
where, $\mathcal{M}$ represents the full sky map, and $\mathcal{F}\in\{{\rm T},{\rm Q},{\rm U}\}$. $\rm Smooth$} represents the telescope’s beam, and we have chosen a simple Gaussian beam. Lenspyx is an abbreviation for the process of converting unlensed maps into lensed maps. $N$ represents the noise component in the image. We also considered the impact of instrumental effects and used a Gaussian beam with a full width at half maximum (FWHM) of $7.9$ arcmins to simplify the processing.
II.1 The CMB temperature and lensing potential
The CMB radiation field is represented by the temperature anisotropy, denoted as T(n^)T($\hat${n}), and the polarization, denoted as P(n^)P($\hat${n}), in the spatial direction $\hat{n}$ of the celestial sphere. We observe CMB temperature changes projected on a 2D spherical surface sky, and it is now habitual in the literature to expand the temperature field using spherical harmonics.
@@ II.1 The CMB temperature and lensing potential
All the information contained in the temperature field T(n^)T($\hat${n}) is included in the space-time dependent amplitudes $a_{\ell m}$, where $a_{\ell m}$ represents the spherical harmonic coefficient and can be expressed using the following formula,
@@
$a_{\ell m}a_{\ell^{\prime}m^{\prime}}^{*}>=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}C_{\ell},$ (4)
@@
We can simulate the angular power spectrum $C_{\ell}$ with CAMB or CLASS Lucca:2019rxf; DiDio:2013bqa, as shown in Fig. 1. The figure demonstrates that as the scale decreases, the difference between lensed and unlensed data amplifies. This phenomenon occurs due to the deflection imparted on CMB photons by each encountered potential, resulting in an accumulation of effects on the CMB power spectrum at small scales. Consequently, the lensing effect distorts the original power spectrum and introduces non-Gaussianity in the lensed CMB Lewis:2006fu. It should be noted that non-linear evolution also contributes to an increase in power on smaller scales.
Weak lensing of the CMB deflects photons coming from an original direction $\hat{n}^{\prime}$ on the last scattering surface to direction $\hat{n}$ on the observed sky, so a lensed CMB temperature field, $T(\theta,\phi)$, is given by $\widetilde{X}(\hat{n})=X(\hat{n}^{\prime})$ in terms of the unlensed field $X=T$ Lewis:2005tp. Thus the position in the sky where we finally see the CMB photons is determined by the integral of the gravitational potential along the line of sight to the last scattering surface.
@@
$),\widetilde{T}(\hat{n})=T(\hat{n}^{\prime})=T(\hat{n}+{{\rm\nabla}}\psi(\hat{n}$)), (6)
[…]
$\psi(\hat{n})=2\int_{0}^{\chi_{\star}}{\rm d}\chi(\frac{\chi_{\star}-\chi}{\chi_{\star}\chi})\Psi(\hat{n},\eta),$ (7)
@@
$C_{\ell}^{\psi\psi}=16\pi\int\frac{{\rm d}k}{k}P_{R}(k)(\int_{0}^{\chi_{\star}}{\rm d}\chi(\frac{\chi_{\star}-\chi}{\chi_{\star}\chi})T(k,\eta)k_{\ell}(k\chi))^{2},$ (8)
where $T(k,\eta)$ is the appropriate transfer function, $P_{R}(k)$ is the primordial power spectrum.
@@
To acquire additional datasets and prepare for subsequent polarization analysis, we conducted experiments using two key cosmological parameters: the scalar amplitude ($A_{s}$) and the tensor-to-scalar ratio ($r$), both of which play crucial roles in determining the temperature and polarization power spectra of the CMB. Specifically, we sampled five values of $A_{s}$ ($2.0\times 10^{-9}$ to $2.2\times 10^{-9}$) and ten values of $r$ linearly spaced between $0.001$ and $0.01$, resulting in a total of 30 CMB TT angular power spectra and corresponding sky maps.
In the top panel of Figure 1, the red solid line and the pink dashed line illustrate the lensed TT angular power spectrum and the unlensed TT power spectrum , with the parameters $A_{s}=2.1\times 10^{-9}$ and $r=0.005$, respectively. The ratio between them is shown with a red dashed line in the bottom panel. On large scales, the anisotropies of the CMB are primarily dominated by emissions from the last scattering surface at a redshift of $z\sim 1100$. However, on smaller scales, the CMB is more significantly influenced by what are known as secondary effects. These secondary anisotropies arise as a result of interactions between CMB photons and matter along the line of sight.
Our work is grounded in the analysis of full-sky maps, thus requiring the use of the Lenspyx tool to transform the generated $30$ angular power spectra into corresponding temperature sky maps. To investigate the impact of lensing potential on small-scale structures, we conducted simulation experiments at a high resolution of $N_{\rm side}=2048$. To ensure that the simulation results more closely match actual observational data, we accounted for the smoothing effect of the telescope beam on signals during the computation process, use a FWHM of $\theta_{{\rm FWHM}}=8.3$ arcmins.
II.2 The CMB polarization
CMB polarization is measured through time-averaged Stokes parameters – measures of linear polarization of the electric field aligned orthogonally to the line of sight’s (LOS) Cartesian axes. Due to the nature of Thomson scattering, Q and U are ample in describing CMB polarization, given their inability to generate circular polarization DODELSON2021373. Astronomical observations reveal that Stoke parameters Q and U are linked by a constant and relative $45{\rm{}^{\circ}}$ degree rotation around LOS – whereas the reference frame may revolve freely around it. Bundling the reference frame with a fixed set of normal vectors in rotation will directly link Q and U to the E and B modes.
If we rotate Q and U by an angle $\alpha$ on the plane that is perpendicular to the direction of $\hat{n}$, we obtain the following solution Zaldarriaga:1996xe; zaldarriaga1998cosmic; Rotti:2018pzi,